1. Calculate the resistance of a copper wire that is 3-meter long and has a diameter of 0.005m.

2. The resistance of a piece of carbon at 200C is 10 Ω. What is its resistance at (a) 300C &(b) 100C?

providing the given, required, formula, solution/substitution of data,


1
Expert's answer
2022-03-28T11:03:11-0400
  1. Given:

l = 3 m,

d = 0.005 m,

"\rho=1.7\cdot 10^{-8} Ohm\cdot m."

Required:

R - ?


Solution:

"R=\rho\frac{l} {S} =\rho\frac{4l} {\pi d^2}=1.7\cdot 10^{-8}\cdot \frac{4\cdot 3} {3.14\cdot 0.005^2}=0.0026\space Ohm."


2.

Given:

"R_{0}= 10\space Ohm,"

"\alpha=-0.0005\space 1/\degree C,"

"t_0=20 \degree C,"

"t_1=30 \degree C,"

"t_2=10 \degree C."

Required:

"R_1" - ?

"R_2" - ?


Solution:

"R_1=R_0(1+\alpha(t_1-t_0))=10(1+(-0.0005)(30-20))=9.95\space Ohm."

"R_2=R_0(1+\alpha(t_2-t_0))=10(1+(-0.0005)(10-20))=10.05\space Ohm."